# An integrated widely tunable linear isolator based on electro-optic Autler-Townes splitting

### v1

@@ Abstract
Optical isolators are essential for laser protection and robust signal routing, but the incorporation of the necessary magneto-optic (MO) materials in foundries has remained a challenge. As an alternative, several integrated non-magnetic isolators based on acousto-optic (AO) and electro-optic (EO) spatio-temporal modulation have been proposed. Unlike MO isolators, these solutions are wavelength agnostic, though few published demonstrations reach performance that is comparable to MO devices. The most significant remaining concerns are on mitigating undesirable sidebands, lowering power consumption, achieving wide bandwidth or wide tunability, and having a design that is practical to deploy. Here we demonstrate a compact EO optical isolator, using thin film lithium niobate, that produces a very high isolation figure of merit ($>32$ dB contrast per dB of insertion loss) with THz-scale (8 nm) tunability and, due to its architecture, achieves linear operation with negligible sideband generation. The device uses a 20x smaller interaction length and consumes 64x less power compared to the state-of-the-art EO and AO architectures, without being subject to many of their technical constraints.
@@
Magneto-optic isolators have persisted as the dominant non-reciprocal technology. However, there remain unsolved challenges with bringing the necessary magneto-optic materials onto integrated photonics platforms [bi_-chip_2011, ghosh_ceyigsilicon--insulator_2012, huang_dynamically_2017, zhang_monolithic_2019, du_monolithic_2018, yan_waveguide-integrated_2020, yan_ultra-broadband_2024], primarily due to foundry constraints, but also due to the requirement of magnetic biasing, and the large absorption and strong chromatic dependence of the magneto-optic materials. In recent years, there have emerged alternative electrically-driven non-reciprocal systems that use either acousto-optic (AO) [hwang_all-fiber-optic_1997, sohn_time-reversal_2018, kittlaus_electrically_2020, sohn_electrically_2021, tian_magnetic-free_2021] or electro-optic (EO) [bhandare_novel_2005, doerr_optical_2011, lira_electrically_2012, tzuang_non-reciprocal_2014, doerr_silicon_2014, dostart_optical_2021, yu_integrated_2023, orsel_electro-optic_2023, gao_thin-film_2024, orsel_giant_2024] symmetry breaking to produce a direction sensitive optical response. The general principle behind these devices is to generate a traveling wave index modulation, or spatio-temporal modulation, from which light experiences some combination of non-reciprocal operations. Non-magnetic isolators that use these approaches have been able to exhibit wide bandwidth non-reciprocal effects [doerr_silicon_2014, yu_integrated_2023, gao_thin-film_2024], very low insertion loss [sohn_electrically_2021, tian_magnetic-free_2021, yu_integrated_2023], ultra-high contrast [bhandare_novel_2005, tian_magnetic-free_2021, sohn_electrically_2021, yu_integrated_2023, gao_thin-film_2024], and linear response [sohn_electrically_2021, tian_magnetic-free_2021], though currently there is no single solution that delivers on all of these metrics simultaneously.
@@
A broad category of non-magnetic isolators leverage non-reciprocal interference induced by spatio-temporal modulation [bhandare_novel_2005, doerr_optical_2011, tzuang_non-reciprocal_2014, dostart_optical_2021, yu_integrated_2023, gao_thin-film_2024, orsel_giant_2024]. These devices are relatively easy to produce, but they require precise tuning of the amplitude and phase of the electrical stimulus for best performance, and also generate a very large number of spurious sidebands. The second broad category of non-magnetic isolators leverage non-reciprocal inter-modal conversion [hwang_all-fiber-optic_1997, lira_electrically_2012, sohn_time-reversal_2018, sohn_electrically_2021, tian_magnetic-free_2021, kittlaus_electrically_2020], also driven by spatio-temporal modulation. These devices do not require precise control of the electrical stimuli, as the non-reciprocity is determined by the availability of specific optical modes and not through interference effects. Additionally, all undesirable sidebands are suppressed due to the absence of suitable optical states. However, these devices require stringent phase matching in both frequency and momentum space which increases the design complexity.
Since isolators can be cascaded, a key figure of merit is the ratio of isolation contrast to insertion loss (IC/IL) for the optical carrier signal. To date, only two reported non-magnetic chip-scale isolators [sohn_electrically_2021, yu_integrated_2023] have reached IC/IL $>30$ dB/dB, i.e. performance that is considered comparable to commercial off-chip magneto-optic devices. The AO isolator demonstrated by Sohn et al. [sohn_electrically_2021] (and the similar design by Tian et al. [tian_magnetic-free_2021]) leverages non-reciprocal inter-modal conversion within a resonator to break the spectral symmetry of its counter-propagating modes. The resonator is then used as a chiral absorber by critical shunt coupling to a waveguide so that, for some range of wavelengths, there is no interaction with light propagating in the forward direction along the waveguide, but simultaneously there is very strong extinction for light propagating in the backward direction. Importantly, all undesirable modulation sidebands are strongly suppressed as they are not phase-matched to any of the available optical states. The forward insertion loss and the residual first-order sideband in the forward direction both diminish monotonically with increasing drive strength, reaching fully linear isolator behavior in the strong-coupling regime with very large drive. However, since the acoustic dispersion locks together the stimulus frequency and momentum, the strict optical-acoustic phase matching is only satisfied between a limited set of optical modes and cannot be extended elsewhere in the spectrum. On the other hand, the EO isolator demonstrated by Yu et al. [yu_integrated_2023] uses traveling wave phase modulation in a long (1.75 cm interaction length) waveguide to produce direction-dependent interference. High transmission can be achieved over a relatively wide bandwidth in the forward direction by tuning the RF drive frequency and length of the modulation region. The RF drive amplitude is then separately fine-tuned to fully deplete any returning signal in the backward direction, but due to the design, the spectral energy is up- and down-scattered into sidebands up to 4th-order. This necessitates the use of a tunable filter which ultimately limits the instantaneous operational bandwidth to that of the filter.
Importantly, both of these device types [sohn_electrically_2021, tian_magnetic-free_2021, yu_integrated_2023] draw considerable RF power. The AO demonstrations [sohn_electrically_2021, tian_magnetic-free_2021] require power to be converted from RF to acoustic domain, and the acoustic waves subsequently propagate away from the device or are dissipated. The EO demonstration [yu_integrated_2023] draws RF power from the signal generator, which propagates through the device, and is subsequently dissipated at a load.
An ideal optical isolator should exhibit ultra-low insertion loss and high directional contrast, with linearity (minimal sidebands or frequency shifts) and linear operation (input-output relation), all with low power consumption. It should also be readily tunable over a wide spectrum. For in-the-field robustness, it is preferable not to rely on free surfaces (e.g. for acoustic reflections) and it should be possible to encapsulate within a cladding for protection. An optimal solution would likely combine the best features of both of the above record-setting approaches [sohn_electrically_2021, yu_integrated_2023].
In this work we experimentally demonstrate a compact electro-optic isolator that exhibits IC/IL $>30$ dB/dB, reaches near-ideal isolator operation with minimal sidebands, and is tunable over a wide wavelength range (we demonstrate this over 1542 – 1550 nm). The design leverages the architectural advantages of the AO-based chiral shunt isolator [sohn_electrically_2021, tian_magnetic-free_2021] while introducing multiple EO innovations that solve phase-matching challenges, enable enhanced EO tuning, and permit cladding of the device without requiring any free surfaces. Notably, the device has a very compact footprint and uses significantly less power than prior state-of-the-art AO and EO isolators.
@@
At the core of this isolator is a two-level “photonic atom” [zhang_electronically_2019] whose optical density of states (DoS) exhibits large chiral asymmetry due to an induced photonic Autler-Townes splitting (p-ATS) effect [peng_what_2014, kim_complete_2017, shi_nonreciprocal_2018, zhang_electronically_2019, sohn_electrically_2021]. This photonic atom is shunt coupled to a bus waveguide to produce isolation (Fig. 1a). A straightforward approach to producing the photonic atom is with a whispering gallery or racetrack resonator that supports only two non-degenerate optical mode families with distinct dispersions. A unidirectional stimulus with non-zero propagating momentum is then applied to the resonator, ensuring that the phase-matching for inter-modal conversion is only satisfied for one direction of propagation (Fig. 1b). When the stimulus-induced coupling rate $G_{\textrm{stim}}>\sqrt{\kappa_{1}\kappa_{2}}$, where $\kappa_{1,2}$ are the loss rates for the individual optical modes, the system enters the strong coupling regime where the non-degenerate optical modes are hybridized and exhibit splitting, analogous to the splitting of electronic states in the conventional ATS phenomenon. Importantly, this p-ATS effect modifies the optical DoS only for the phase-matched direction [sohn_electrically_2021, kim_complete_2017, shi_nonreciprocal_2018]. For the counter-propagating direction, the original optical DoS remains unmodified. The previous AO implementations of this concept used single input propagating acoustic wave [sohn_electrically_2021] and 3-input acoustic standing wave [tian_magnetic-free_2021] stimuli to accomplish the phase-matching. Here we aim to replace this with an EO stimulus to facilitate tuning and encapsulation. However, the major challenge is that, unlike acoustic phonons which possess substantial momentum [sohn_electrically_2021, kittlaus_electrically_2020], RF traveling waves have very high phase velocity and are unable to bridge the required momentum gaps at practical frequencies.
@@
Figure 2: Design and implementation of the electro-optic p-ATS isolator. (a) The cross-sectional schematic of the double-racetrack resonator shows the lithium niobate (LN) waveguide structure (dark blue), electrode configuration (yellow/orange), and oxide cladding (grey). (b) Finite element simulation of the TEeven and TEodd optical modes supported by the resonator. (c) Finite element simulation of the RF mode showing the transverse mirror symmetric electric-field profile. (d) A plan-view schematic of the 3-phase EO stimulus section is presented. The split-electrode structure is driven through three RF bus lines with $120^{\circ}$ relative phase shift with matched amplitudes at driving frequency $\Omega$. (e) A top-view optical microscope image of the fabricated EO isolator. Upper section of the racetrack resonator shows the 3-phase EO modulator with the inset showing a zoomed-in view of a single electrode. Lower section of the resonator enables EO fine-tuning of the isolator.
@@
Our solution is to invoke a periodic EO modulator that generates a synthetic stimulus momentum [orsel_electro-optic_2023]. This approach also alleviates the phase-matching challenge across a wide spectrum by making the frequency a “free choice” variable that can be selected appropriately during operation. The schematic for our specific modulator is shown later in Fig. 2d, where a controlled relative phase shift $\Delta\phi=120^{\circ}$ is applied between subsequent electrodes having pitch $L$. This configuration establishes a dominant spatial wavenumber of $q=\Delta\phi/L$. The stimulus momentum is therefore set by an initial lithographic choice of $L$, while the stimulus frequency $\Omega$ is freely selectable during operation.
@@
The next major requirement is to ensure a fixed azimuthal momentum separation between the optical mode pairs over a wide spectral range. The previous AO designs [sohn_electrically_2021, tian_magnetic-free_2021] required precise dimension control of the resonator to obtain the required momentum separation between the modes, but this causes the device yield to be very sensitive to small fabrication variations. Moreover, since the optical dispersions are different, the phase matching in a selected device is only achievable between a limited set of optical mode pairs, which restricts dynamic tuning of the isolation wavelength. Accordingly, our second key innovation is to instead use a double-racetrack resonator [xu_mutual-coupling_2022, zhang_broadband_2022], composed of identical component waveguides having identical underlying dispersion (Fig. 2a). This allows us to very easily control the momentum separation of the TEeven and TEodd optical modes (Fig. 2b) across a very wide spectrum, even though the frequency separation can vary slightly (as seen later in Fig. 5a). The combination of the above two design features – the freedom to choose the stimulus frequency, and the very steady momentum separation – enables the isolation effect to be coarsely tuned to optical mode pairs across a very wide spectral range.
@@
where $\mathcal{E}^{\textrm{even}}({r_{\bot}})$, ${\mathcal{E}}^{\textrm{odd}}({r_{\bot}})$, and ${\mathcal{E}}^{\textrm{RF}}(r_{\bot})$ are the transverse field distribution of optical and RF modes respectively (Fig. 2b,c) and the subscripts $ijk$ refer to coordinate axes for the electro-optic tensor $\mathfrak{r}_{ijk}$ (the mathematical details are presented in [orsel_electro-optic_2023]). For our specific device, the best RF electrode configuration requires a two-layer structure. The lower layer supports the uninterrupted ground electrodes while the upper layer is split into the multi-phase stimulus electrode (Fig. 2a,d).
@@ Results
We fabricated this optical isolator (Fig. 2e) using a 600 nm X-cut thin-film lithium niobate on insulator (TFLN) integrated photonics platform with an oxide cladding and two-layer gold electrodes to generate the required electro-optic stimuli. The fabrication process and key geometric features are described in Methods. Lithium niobate is well-suited for electro-optics due to its large EO coefficient ($r_{33}\approx 30$ pm/V) and wide bandgap, offering a broad transparency range of 250 to 5300 nm with propagation losses below 0.1 dB/cm [weis_lithium_1985, desiatov_ultra-low-loss_2019]. The resonator design targets an optical momentum gap of 13000 – 15000 rad/m. Accordingly, two periods of the 3-phase RF stimulus electrodes were patterned on one side of the resonator, and their pitch was selected at $L=150$ µm, resulting in a total interaction length of 900 µm. The stimulus electrode cross-section (Fig. 2a) was designed to generate a mirror-symmetric transverse electric field profile along the LN Z-axis (Fig. 2c) to efficiently harness the $r_{33}$ coefficient for modulation. We additionally fabricated a set of EO tuning electrodes on the opposite side of the resonator.
@@
Figure 3: Experimental demonstration and characterization of isolation. (a) Transmission measured through the waveguide shows the two optical mode families hosted by the resonator. We chose the optical mode pair marked by $\color[rgb]{1,0,0}\bm{\star}$ located near 193.98 THz (1545.5 nm) to conduct our experiments. (b) With RF stimulus set to $\Omega=4.8$ GHz, which matches the modal frequency gap, we measured evolution of the through-waveguide transmission for the $\color[rgb]{1,0,0}\bm{\star}$ mode pair. As the applied RF voltage (VRF) is increased, the p-ATS phenomenon appears for forward light propagation direction while the backward direction remains strongly absorbing. (c) We present a detailed view of the superimposed forward and backward measurements for the 15.88 V case, corresponding to a measured $G_{eo}$ of approximately 0.89 GHz.

### v2

# An integrated multi-THz tunable linear isolator based on electro-optic non-reciprocal strong coupling

@@ Abstract
Optical isolators are essential for laser protection and robust signal routing, but the incorporation of the necessary magneto-optic (MO) materials in foundries has remained a challenge. As an alternative, several integrated non-magnetic isolators based on acousto-optic (AO) and electro-optic (EO) spatio-temporal modulation have been proposed. Unlike MO isolators, these solutions are wavelength agnostic, though few published demonstrations reach performance that is comparable to MO devices. The most significant remaining concerns are on mitigating undesirable sidebands, achieving wide bandwidth or wide tunability, and having a design that is practical to deploy. Most of these challenges can be addressed through non-reciprocal strong coupling between waveguide or resonator modes, with the intent to produce extremely asymmetric optical dispersion, but this has never been accomplished with electro-optics. Here we demonstrate a compact EO optical isolator, using thin film lithium niobate, that is the first EO device to reach the non-reciprocal strong coupling regime. In this new regime, the isolator produces a very high isolation figure of merit (>32>32 dB contrast per dB of insertion loss, 47.7 dB isolation contrast with 1.45 dB insertion loss) and, due to its architecture, achieves linear operation with negligible sideband generation. We additionally demonstrate THz-scale (8 nm) tunability of the isolation band that is not fundamentally limited, and can be extended to multi-THz operation.
@@
Magneto-optic isolators have persisted as the dominant non-reciprocal technology. However, there remain unsolved challenges with bringing the necessary magneto-optic materials onto integrated photonics platforms [bi_-chip_2011, ghosh_ceyigsilicon--insulator_2012, huang_dynamically_2017, zhang_monolithic_2019, du_monolithic_2018, yan_waveguide-integrated_2020, yan_ultra-broadband_2024], primarily due to foundry constraints, but also due to the requirement of magnetic biasing, and the large absorption and strong chromatic dependence of the magneto-optic materials. In recent years, there have emerged alternative electrically-driven non-reciprocal systems that use either acousto-optic (AO) [hwang_all-fiber-optic_1997, sohn_time-reversal_2018, kittlaus_electrically_2020, sohn_electrically_2021, tian_magnetic-free_2021, cheng_terahertz-bandwidth_2025] or electro-optic (EO) [bhandare_novel_2005, doerr_optical_2011, lira_electrically_2012, tzuang_non-reciprocal_2014, doerr_silicon_2014, dostart_optical_2021, yu_integrated_2023, orsel_electro-optic_2023, gao_thin-film_2024, orsel_giant_2025] symmetry breaking to produce a direction sensitive optical response. The general principle behind these devices is to generate a traveling wave index modulation, or spatio-temporal modulation, from which light experiences some combination of non-reciprocal operations. Non-magnetic isolators that use these approaches (Supplementary Data Table 1) have been able to exhibit wide bandwidth non-reciprocal effects [doerr_silicon_2014, yu_integrated_2023, gao_thin-film_2024, cheng_terahertz-bandwidth_2025], very low insertion loss [sohn_electrically_2021, tian_magnetic-free_2021, yu_integrated_2023], ultra-high contrast [bhandare_novel_2005, tian_magnetic-free_2021, sohn_electrically_2021, yu_integrated_2023, gao_thin-film_2024], and linear response [sohn_electrically_2021, tian_magnetic-free_2021, cheng_terahertz-bandwidth_2025], though currently there is no single solution that delivers on all of these metrics simultaneously.
@@
A broad category of non-magnetic isolators leverage non-reciprocal interference induced by spatio-temporal modulation [bhandare_novel_2005, doerr_optical_2011, tzuang_non-reciprocal_2014, dostart_optical_2021, yu_integrated_2023, gao_thin-film_2024, orsel_giant_2025]. These devices are relatively easy to produce, but they require precise tuning of the amplitude and phase of the electrical stimulus for best performance, and also generate a very large number of spurious sidebands. The second broad category of non-magnetic isolators leverage non-reciprocal inter-modal conversion [hwang_all-fiber-optic_1997, lira_electrically_2012, sohn_time-reversal_2018, sohn_electrically_2021, tian_magnetic-free_2021, kittlaus_electrically_2020, cheng_terahertz-bandwidth_2025], also driven by spatio-temporal modulation. These devices do not require precise control of the electrical stimuli, as the non-reciprocity is determined by the availability of specific optical modes and not through interference effects. Additionally, all undesirable sidebands are suppressed due to the absence of suitable optical states. However, these devices require stringent phase matching in both frequency and momentum space which increases the design complexity.
A particularly effective version of the inter-modal approach is the chiral shunt design demonstrated by Sohn et al. [sohn_electrically_2021] and the similar design by Tian et al. [tian_magnetic-free_2021], which can simultaneously achieve ultra-low insertion loss and large directional contrast, with linearity (minimal frequency shifted sidebands) and linear operation (input-output relation). However, since this design relies on AO inter-modal coupling, the phase-matching is very strict and the isolation is both narrowband and difficult to spectrally tune. This is because the acoustic dispersion locks the frequency and momentum of the acoustic stimulus, and because acoustic transducers provide only a limited driving bandwidth. These AO designs also require power to be converted from an input RF electrical stimulus into acoustic domain, and the acoustic waves subsequently propagate away from the device or are dissipated.
In this work we experimentally demonstrate a compact EO isolator that brings transformative improvements to the chiral shunt design – significantly relaxing the phase matching, achieving multi-THz tunability, and enabling protective oxide cladding. We demonstrate tuning of this isolator over a wide 8 nm span, that is only limited by our experimental capability, and we confirm that it can consistently maintain very large isolation contrast (IC >41>41 dB) with very low insertion loss (IL <2<2 dB) over this tuning range. Our best result shows IC =47.7=47.7 dB simultaneously with IL =1.45=1.45 dB. In doing so, we also achieve the first EO device to operate in the non-reciprocal strong coupling regime. There are three major design innovations that enable this result. First, RF excitations have high phase velocity and do not naturally carry sufficient momentum to allow for the required phase-matching, which we resolve here with a synthetic momentum approach. Second, we simplify the optical mode placement in frequency-momentum space by means of a double racetrack resonator that allows for flexible momentum spacing by design. Finally, we implement an EO electrode set that can successfully break the orthogonality between these optical modes while achieving a large coupling rate.
@@
At the core of this isolator is a two-level “photonic atom” [zhang_electronically_2019] whose optical density of states (DoS) exhibits large chiral asymmetry due to an induced photonic Autler-Townes splitting (p-ATS) effect [peng_what_2014, kim_complete_2017, shi_nonreciprocal_2018, zhang_electronically_2019, sohn_electrically_2021]. This photonic atom is shunt coupled to a bus waveguide to produce isolation (Fig. 1a). A straightforward approach to producing the photonic atom is with a whispering gallery or racetrack resonator that supports only two non-degenerate optical mode families with distinct dispersions. A unidirectional stimulus with non-zero propagating momentum is then applied to the resonator, ensuring that the phase-matching for inter-modal conversion is only satisfied for one direction of propagation (Fig. 1b). When the stimulus-induced coupling rate $G_{\textrm{stim}}>\sqrt{\kappa_{1}\kappa_{2}}$, where $\kappa_{1,2}$ are the loss rates for the individual optical modes, the system enters the strong coupling regime where the non-degenerate optical modes are hybridized and exhibit splitting, analogous to the splitting of electronic states in the conventional ATS phenomenon (a detailed coupled mode theory treatment can be found in the Supplementary Materials in [sohn_electrically_2021]). Importantly, this mode splitting effect modifies the optical DoS only for the phase-matched direction [sohn_electrically_2021, kim_complete_2017, shi_nonreciprocal_2018]. For the counter-propagating direction, the original optical DoS remains unmodified. The resonator is then used as a chiral absorber by critical shunt coupling to a waveguide so that, for some range of wavelengths, there is no interaction with light propagating in the forward direction along the waveguide, but simultaneously there is very strong extinction for light propagating in the backward direction. Importantly, all undesirable modulation sidebands are strongly suppressed as they are not phase-matched to any of the available optical states. Coupled mode theory for ATS type isolators [sohn_electrically_2021] predicts that forward insertion loss and the residual first-order sideband in the forward direction both diminish monotonically with increasing drive strength, reaching fully linear isolator behavior as $G_{\textrm{stim}}\rightarrow\infty$.
The previous AO implementations of this concept used single input propagating acoustic wave [sohn_electrically_2021] and 3-input acoustic standing wave [tian_magnetic-free_2021] stimuli to accomplish the phase-matching. Here we aim to replace this with an EO stimulus to facilitate improved tunability and robust cladding. However, unlike acoustic phonons which possess substantial momentum [sohn_electrically_2021, kittlaus_electrically_2020], RF traveling waves have very high phase velocity and are unable to bridge the required momentum gaps at practical frequencies. To date, there has been no demonstration of non-reciprocal strong coupling in an integrated electro-optic device – in either waveguide or resonator geometry – which can be potentially transformative for non-reciprocal photonics technologies.
@@
Figure 2: Design and implementation of the electro-optic isolator. (a) The cross-sectional schematic of the double-racetrack resonator shows the lithium niobate (LN) waveguide structure (dark blue), electrode configuration (yellow/orange), and oxide cladding (grey). (b) Finite element simulation of the TEeven and TEodd optical modes supported by the resonator. (c) Finite element simulation of the RF mode showing the transverse mirror symmetric electric-field profile. (d) A plan-view schematic of the 3-phase EO stimulus section is presented. The split-electrode structure is driven through three RF bus lines with $120^{\circ}$ relative phase shift with matched amplitudes at driving frequency $\Omega$. (e) A top-view optical microscope image of the fabricated EO isolator. Upper section of the racetrack resonator shows the 3-phase EO modulator with the inset showing a zoomed-in view of a single electrode. Lower section of the resonator enables EO fine-tuning of the isolator.
@@
Our solution is to invoke a periodic EO modulator that generates a synthetic stimulus momentum [orsel_electro-optic_2023]. This approach also alleviates the phase-matching challenge across a wide spectrum by making the frequency a “free choice” variable that can be selected appropriately during operation. The schematic for our specific modulator is shown later in Fig. 2d, where a controlled relative phase shift $\Delta\phi=120^{\circ}$ is applied between subsequent electrodes having pitch $L$. This configuration establishes a dominant spatial wavenumber of $q=\Delta\phi/L$ (harmonics are also generated but are rejected by the resonator). The stimulus momentum is therefore set by an initial lithographic choice of $L$, while the stimulus frequency $\Omega$ is freely selectable during operation. This synthetic momentum $q$ directly replaces the acoustic momentum from the AO designs [sohn_electrically_2021, tian_magnetic-free_2021].
@@
The next major requirement is to ensure a fixed azimuthal momentum separation between the optical mode pairs over a wide spectral range. The previous AO designs [sohn_electrically_2021, tian_magnetic-free_2021] required precise dimension control of the resonator to obtain the required momentum separation between the modes, but this causes the device yield to be very sensitive to small fabrication variations. Moreover, since the optical dispersions are different, the phase matching in a selected device is only achievable between a limited set of optical mode pairs, which restricts dynamic tuning of the isolation wavelength. Accordingly, our second key innovation is to instead use a double-racetrack resonator [xu_mutual-coupling_2022, zhang_broadband_2022], composed of identical component waveguides having identical underlying dispersion (Fig. 2a). This allows us to very easily control the momentum separation of the TEeven and TEodd optical modes (Fig. 2b) across a very wide spectrum, even though the frequency separation can vary slightly (as seen later in Fig. 5). The combination of the above two design features – the freedom to choose the stimulus frequency, and the very steady momentum separation – enables the isolation effect to be coarsely tuned to optical mode pairs across a very wide spectral range.
@@
where $\mathcal{E}^{\textrm{even}}({r_{\bot}})$, ${\mathcal{E}}^{\textrm{odd}}({r_{\bot}})$, and ${\mathcal{E}}^{\textrm{RF}}(r_{\bot})$ are the transverse field distribution of optical and RF modes respectively (Fig. 2b,c) and the subscripts ijkijk refer to coordinate axes for the electro-optic tensor $\mathfrak{r}_{ijk}$ (the mathematical details are presented in [orsel_electro-optic_2023]). For our specific device, the best RF electrode configuration requires a two-layer structure. The lower layer supports the uninterrupted ground electrodes while the upper layer is split into the multi-phase stimulus electrode (Fig. 2a,d).
@@ Results
We fabricated this optical isolator (Fig. 2e) using a 600 nm X-cut thin-film lithium niobate on insulator (TFLN) integrated photonics platform with an oxide cladding and two-layer gold electrodes to generate the required electro-optic stimuli. The fabrication process and key geometric features are described in Methods. Lithium niobate is well-suited for electro-optics due to its large EO coefficient ($r_{33}\approx 30$ pm/V) and wide bandgap, offering a broad transparency range of 250 to 5300 nm with propagation losses below 0.1 dB/cm [weis_lithium_1985, desiatov_ultra-low-loss_2019]. The resonator design targets an optical momentum gap of 13000 – 15000 rad/m corresponding to an azimuthal order difference of $\Delta M_{\textrm{optical}}=7$. Accordingly, two periods of the 3-phase RF stimulus electrodes were patterned on one side of the resonator, and their pitch was selected at $L=150$ µm, resulting in a total interaction length of 900 µm (stimulus wavelength is then 450 µm, which matches $M_{\textrm{stimulus}}=7$ around the racetrack). The stimulus electrode cross-section (Fig. 2a) was designed to generate a mirror-symmetric transverse electric field profile along the LN Z-axis (Fig. 2c) to efficiently harness the $r_{33}$ coefficient for modulation. We additionally fabricated a set of EO tuning electrodes on the opposite side of the resonator.
@@
[…]
{adjustwidth*}
-1in-1in
Figure 4: Sideband analysis during isolator operation. (a) The measured first-order sidebands as a function of optical detuning, normalized to the input optical carrier, for the case shown in Fig. 3c. (b) These beat spectra obtained from the heterodyne detection system reveal the carrier and its sidebands (see Supplementary Data Fig. 1), normalized to the input optical carrier. Measurements are performed with a $\Omega_{r}=100$ MHz frequency shift applied to the reference arm of the interferometer; thus $\Omega_{r}$ measures the carrier, $\Omega\pm\Omega_{r}$ correspond to the first-order Stokes and anti-Stokes sideband respectively, and $2\Omega\pm\Omega_{r}$ correspond to the second-order Stokes and anti-Stokes sideband respectively. The anti-Stokes sidebands are notably absent, and beat notes at $\Omega,2\Omega$ correspond to parasitics in the measurement system. (c) We plot the normalized carrier transmission and the Stokes sideband power in forward direction, and backward carrier transmission, all normalized to input optical carrier, as a function of RF voltage at $\Delta$ = 0. The solid line is a prediction from coupled mode theory [sohn_time-reversal_2018, sohn_electrically_2021] using the experimentally measured device parameters.
Demonstration of electro-optic isolation – Fig. 3 presents an example experiment on the isolator. First, we measured optical transmission through the waveguide in the telecom band, around 1550 nm, which reveals two distinct optical mode families of the resonator (Fig. 3a) with separation of a few GHz. The modes can be identified as TEeven and TEodd based on both coupling to the waveguide and their dispersive effect supported through finite element simulations. Next, we experimentally measure the through-waveguide transmission using a heterodyne detection system (see Methods) which enables resolved measurements on the carrier transmission, as well as any sideband signals that are generated. In this experiment we chose an optical mode pair near 193.98 THz (1545.50 nm) as identified in Fig. 3a with frequency gap of approximately 4.8 GHz and individual loss rates $\kappa_{\textrm{even}}=0.28$ GHz and $\kappa_{\textrm{odd}}=$ 0.53 GHz. The TEodd mode is critically coupled and provides large extinction for the waveguide transmission, while the TEeven mode is more inward located and is therefore undercoupled. We accordingly apply a 3-phase RF stimulus at $\Omega=4.8$ GHz to the electrodes. We also note that this RF drive frequency does not have to be precisely matched to the frequency gap and the isolator continues to operate as long as the system of modes remains in the strong coupling regime [sohn_electrically_2021]. Fig. 3b presents the evolution of the optical transmission through waveguide as a function of RF voltage (i.e. sinusoidal amplitude of RF stimulus, VRF) for both forward and backward directions. Here, we arbitrarily set the TEodd mode as the zero detuning reference point ($\Delta=0$) for the optical frequency axis. In the absence of any applied RF, the transmission function remains identical in both directions. However, with increasing RF voltage, the p-ATS modal hybridization appears for one direction of propagation only. In Fig. 3c we present a superimposed result at the highest applied RF amplitude of 15.88 V, reaching a peak isolation contrast of 47.7 dB with a forward insertion loss of 1.45 dB. This implies IC/IL ratio of $\approx 32.9$ dB/dB which makes the performance comparable to commercial magneto-optic isolators. By fitting to coupled mode theory [sohn_electrically_2021], we determine that this driving condition corresponds to an EO coupling rate $G_{eo}\approx 0.89$ GHz. We measure a 10 dB IC bandwidth of $\approx 150$ MHz and 20 dB IC bandwidth of $\approx 40$ MHz. We tested this device with an estimated in-waveguide optical power up to 5 mW.
While the EO stimulus on this device has an overall traveling wave characteristic, the propagating momentum is generated only synthetically. We confirm this using a RF S11 reflection measurement at the stimulus electrodes (Supplementary Data Figure 2) which confirms their uniformly high reflection and open circuit (capacitive) behavior [orsel_electro-optic_2023]. The residual loss arises from propagation in the measurement cables. For the specific experiment in Fig. 3, the bench-top signal generator power was set at 28 dBm (630 mW). Since the load at the device is a capacitor (i.e. open circuit), and since the source, cable, and RF probe impedances are all 50 ohms, this corresponds to a sinusoidal voltage on each electrode having amplitude $V_{\textrm{RF}}=2\times\sqrt{2\cdot 50\cdot P_{drive}}=15.88$ V [orsel_electro-optic_2023]. The RF S11 measurement at 4.8 GHz informs us of a fundamental upper limit of only 12.9 mW (11.09 dBm) power consumed by the device. With improved RF design (e.g. impedance matching circuits if using external signal generators) or a custom RF voltage generator on-chip, the power consumption could be reduced significantly. Thus, essentially all of the signal generator voltage (15.88 V) appears directly on the stimulus electrodes due to the open circuit.
Sideband analysis – During experiments we also monitored the anti-Stokes and Stokes sidebands since they are a concern in non-reciprocal devices that use strong spatio-temporal modulation [doerr_optical_2011, doerr_silicon_2014, yu_integrated_2023, gao_thin-film_2024]. Fig. 4a presents measurements of the first-order sidebands as a function of optical detuning $\Delta$, under the operational conditions of Fig. 3c. Naturally, the sideband generation in the forward (phase-matched) direction is shaped by the optical DoS as modified by the mode splitting. In the limit of very strong coupling, the hybrid states will be fully resolved [kim_complete_2017], and at $\Delta=0$ the optical DoS will reach zero causing the forward direction first-order sideband to be eliminated. In the backward (not phase-matched) direction the sidebands are significantly smaller.
Fig. 4b analyzes the higher-order sidebands in the optical spectrum using the heterodyne interferometer (see Methods and Supplementary Data Figure 1) with the input laser centered on the TEodd mode ($\Delta=0$) and reference arm frequency shift of $\Omega_{r}=100$ MHz. We expect that only first-order Stokes scattering will be notable since only scattering to TEeven can take place. The interferometer spectrum shows the carrier (at $\Omega_{r}$), first Stokes (at $\Omega+\Omega_{r}$), and a much smaller second-order Stokes (at $2\Omega+\Omega_{r}$) signal. We normalize the forward and backward measurements relative to the carrier transmission in the forward direction. These measurements confirm that only the first Stokes sideband is of significance at approximately -17.8 dB relative to the carrier. The first anti-Stokes ($\Omega-\Omega_{r}$) and second anti-Stokes ($2\Omega-\Omega_{r}$) are absent, or below the noise floor, as there are no optical states available within the resonator.
These sideband generation behaviors are readily predictable from coupled mode theory [sohn_electrically_2021, kim_complete_2017] and two regimes of operation have been identified. When the applied stimulus $G_{eo}$ is small, the device acts as a non-reciprocal frequency shifting modulator [sohn_time-reversal_2018]. However, once $G_{eo}>\sqrt{\kappa_{\textrm{even}}\cdot\kappa_{\textrm{odd}}}$ the device enters the strong coupling regime where linear isolation becomes available. With increasing drive level the hybrid states get resolved further and the ideality of the device monotonically improves as sideband generation reduces monotonically. This behavior and the close match between experiments and theory is confirmed in Fig. 4c. Importantly, if the auxiliary mode (here the TEeven mode) is fully decoupled from the bus waveguide, any residual scattered Stokes sideband can be completely eliminated.